Name: MAT 145: Topics In Contemporary Math

Operations in Bases
Find each of the following. Be careful which base the work is done in!
1) 11001110(2) + 100011111(2)

Found 2 solutions by Solver92311, MathLover1:
Answer by Solver92311(821)



    11001110
 + 100011111
------------
   111111101



John

My calculator said it, I believe it, that settles it

From
I > Ø

Answer by MathLover1(20850)

there are four rules to followed when adding two binary numbers:
0+0=0
1+0=1
1+1=10 (binary 2)->write 0 and cary 1
1+1+1=11 (binary 3)->write 1 and cary 1
subtracting two binary numbers:
0-0=0
1-0=1
1-1=0
0-1=1 with a borrow of 1

1) 11001110(2) + 100011111(2)
Binary value:
Solution is
...................................1........ 1........ 1........ 1(cary)
0........ 1........ 1........ 0........ 0........ 1........ 1........ 1........ 0
1........ 0........ 0........ 0........ 1........ 1........ 1........ 1........ 1
_______________________________________________________
1........ 1........ 1........ 1........ 0........ 1........ 1........ 0........ 1



11001110 (2)+ 1100011111(2)= 111101101(2)